OpenStudy (anonymous):

Which of the following best defines 3/2^3 Square root of 9 Cube root of 9 Cube root of 3 Square root of 3

OpenStudy (anonymous):

Can you use parentheses to make the problem clearer?

OpenStudy (anonymous):

I'll post a picture one second @pabloecortez07

OpenStudy (anonymous):



OpenStudy (anonymous):

is that okay?

OpenStudy (anonymous):

Not really, you are supposed to raise something to a power, that's just a mixed fraction. Do you mean \[3^\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

yes i mean that, but thats the question on my homework :/ i don't get it really

OpenStudy (anonymous):

Basically what you're doing is taking the cubed root of 9

OpenStudy (anonymous):

so that is the answer?

OpenStudy (anonymous):

\[3^{\frac{ 2 }{ 3 }} = \sqrt[3]{3^{2}}\]

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

@lonelyflowers hey, it'd be super helpful if you could check out hwhelpguy.weebly.com it's a project I started that I hope to expand, thank you :D